OpenStudy (anonymous):

asking to use the implicit differentiation to find dy/dx of 2xy-y^2=1

OpenStudy (anonymous):

first: 2y+2xy'-2yy'=0 Next, solve for y': y' = -2y/(2x-2y)

OpenStudy (anonymous):

the answer should be y/y-x .. can u explain step by step if you get this answer.

OpenStudy (anonymous):

yes, that is just simplifying -2y/(2x-2y)

OpenStudy (anonymous):

factor a -2 out of the numerator and denominator: -2(y)/[-2(y-x)]=y/(y-x)

OpenStudy (anonymous):

thank jjanelle.. ur great